OpenStudy (anonymous):

simplify the expression 3^7 over 3^5

OpenStudy (anonymous):



OpenStudy (anonymous):

Subtract the exponents. 3^7 - 3^5 = 3^(7-5) = 3^2

OpenStudy (anonymous):

it still doesnt match up to the choices i have i dont know what to do this question is worth 5 points but nothing matches up

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

3^35,3^12,139 or 9

OpenStudy (anonymous):

it mut just be a bad question because they isted the question as simplify the expression 3735

OpenStudy (jdoe0001):

hmmm I see now where the "3735" came from :/

OpenStudy (anonymous):

Well 3^2 is 9 so 9 could be the answer...

OpenStudy (jdoe0001):

how about just taking a quick screenshot of the material?

OpenStudy (anonymous):

i can just draw the question out like it is written but my computer doesnt want me to let me take a screenhot for some reason

OpenStudy (anonymous):